OpenStudy (anonymous):

Describe the vertical asymptote(s) and hole(s) for the graph (x-7)(x-7)/(x-7)(x+9)

OpenStudy (anonymous):

\[\frac{(x-7)(x-7)}{(x-7)(x+9)}=\frac{x-7}{x+9}\] "hole" is at x = 7, vertical asymptote is x = -9

OpenStudy (anonymous):

To find vertical asymptotes you cancel what you can (in this case x-7) then set the denominator to zero. The holes are what canceled which was the x-7 set equal to zero.